OpenStudy (love_333):

Use the binomial theorem to expands the binomial. (d-4b)^3

OpenStudy (love_333):

Anybody? Help please.

OpenStudy (holsteremission):

"explains"? Do you mean "expand"? The binomial theorem says \[(x+y)^n=\sum_{k=0}^n\binom nkx^{n-k}y^k\]where \(\dbinom nk=\dfrac{n!}{k!(n-k)!}\). In your example, you can set \(x=d\), \(y=-4b\), and \(n=3\). Then \[\begin{align*} (d-4b)^3&=\sum_{k=0}^3\binom 3kd^{3-k}(-4b)^k\\[1ex] &=\binom 30d^3(-4b)^0+\binom 31d^2(-4b)^1\\ &\qquad+\binom 32d^1(-4b)^2+\binom 33d^0(-4b)^3\\[1ex] &=\cdots \end{align*}\]

OpenStudy (love_333):

@HolsterEmission Yes thats what I meant, Spellcheck must have changed it before I posted the question. I still don't get it? What's the answer?

OpenStudy (love_333):

I don't understand this at all

OpenStudy (love_333):

@sunnnystrong Do you happen to know anything about this?

OpenStudy (sunnnystrong):

Hmm.. not really sure. I'm pretty sure you just expand it? @holsteremission explained it pretty well.

OpenStudy (love_333):

I don't really know anything about this stuff, I just started it so what @holsteremission said didn't really make any sense to me

zepdrix (zepdrix):

Hey Love! It's a `bi`nomial because you have `two` terms inside of the brackets. When you expand a binomial, the powers `count down on the first term`, and `count up on the second term`. \(\rm (x+y)^5=\text{__ }x^5y^0+\text{__ }x^4y^1+\text{__ }x^3y^2+\text{__ }x^2y^3+\text{__ }x^1y^4+\text{__ }x^0y^5\) Each term in the expansion gets a coefficient. If the formula for binomial coefficient is confusing, you can instead look to Pascal's Triangle. This is what the 5th row looks like 1 5 10 10 5 1 Those will be our coefficients,\[\rm (x+y)^5=1x^5y^0+5x^4y^1+10x^3y^2+10x^2y^3+5x^1y^4+1x^0y^5\]

OpenStudy (love_333):

I feel kinda of stupid, I just switched to a new school and this is what they are studying and i've never even seen this stuff before @sunnnystrong

zepdrix (zepdrix):

Try to understand what I posted. It's really just a pattern to follow.

OpenStudy (love_333):

@zepdrix Okay I think I understand the pattern

OpenStudy (sunnnystrong):

@Love_333 you're not stupid haha XD there are also online youtube videos that can help. or khanacademy (just use google)

zepdrix (zepdrix):

Think of your binomial like this,\[\large\rm (d+-4b)^3\]Treat the subtraction as if it's part of the second term.

zepdrix (zepdrix):

So we're going to have a bunch of terms added together after we expand,\[\large\rm =\text{__ }d^3(-4b)+\text{__ }d^2(-4b)+\text{__ }d^1(-4b)+\text{__ }d^0(-4b)\]The first term, d, is counting down from 3.

zepdrix (zepdrix):

And the second term (-4b) is counting up from 0,\[\large\rm =\text{__ }d^3(-4b)^0+\text{__ }d^2(-4b)^1+\text{__ }d^1(-4b)^2+\text{__ }d^0(-4b)^3\]

zepdrix (zepdrix):

And then if we look at Pascal's Triangle, the third row (counting the top row as the 0th row) will give us our coefficients. 1 3 3 1

zepdrix (zepdrix):

\[\large\rm =1d^3(-4b)^0+3d^2(-4b)^1+3d^1(-4b)^2+1d^0(-4b)^3\]

zepdrix (zepdrix):

So we've got the weird pattern stuff out of the way. Now it's just a matter of simplifying.

zepdrix (zepdrix):

Start by replacing the 0 power terms with 1's. (Anything to the zero power is 1).\[\large\rm =1d^3(1)+3d^2(-4b)^1+3d^1(-4b)^2+1(1)(-4b)^3\] Then suppress any 1's that you see. 1 exponents don't mean anything. And also 1 multiplication doesn't mean anything significant.\[\large\rm =d^3+3d^2(-4b)+3d(-4b)^2+(-4b)^3\]It's starting to look a little better, ya?

OpenStudy (love_333):

Yeah, I think so. So whenever you have an equation like this if there are ones do you always take them out like that?

zepdrix (zepdrix):

Yes, one is a stupid number. Always get rid of them if you can :)

OpenStudy (love_333):

Okay, thats good to know!

zepdrix (zepdrix):

From here, we need to expand any powers of (-4b). Recall that when you apply an exponent to a group, you apply it to each thing in the group, \(\large\rm (ab)^x=a^xb^x\)

zepdrix (zepdrix):

So then, \(\large\rm (-4b)^2=(-4)^2b^2\) We'll have to do this with the last term as well.

OpenStudy (love_333):

Okay, so what's the next step after that?

zepdrix (zepdrix):

Simplify the stuff involving the 4's. (-4)^2 = ? (-4)^3 = ?

OpenStudy (love_333):

16 and -64 ?

zepdrix (zepdrix):

\[\large\rm =d^3+3d^2(-4b)+\color{orangered}{3d(-4b)^2+(-4b)^3}\]Ya, so apply the power expansion I was talking about before,\[\large\rm =d^3+3d^2(-4b)+\color{orangered}{3d(-4)^2b^2+(-4)^3b^3}\]and put in those values,\[\large\rm =d^3+3d^2(-4b)+\color{orangered}{3d(16)2b^2+-64b^3}\]

zepdrix (zepdrix):

Recall that multiplication is `commutative`, this means that we can multiply things in any order. So instead of having 3d^2(-4b), we would like to think of it like this, -4*3d^2b bringing all of the number stuff to the front.

zepdrix (zepdrix):

You want to do that with all of them.

zepdrix (zepdrix):

So our second term would look like this after multiplying the number stuff,\[\large\rm =d^3-12d^2b+3d(16)2b^2+-64b^3\]with a -12 in front. Do that with the other terms as well.

OpenStudy (love_333):

Okay so.... \[d^3 = 12d^2b + 48d^2b - 64d^3\] ? Did I do that right?

OpenStudy (love_333):

I feel like I didnt XD

zepdrix (zepdrix):

You accidentally put an equals sign in place of the subtraction sign. Looks good besides that though! Yayyy good job \c:/

OpenStudy (love_333):

I know I just saw that! Really??

zepdrix (zepdrix):

Oh oh and also that third term should be db^2 not d^2b

zepdrix (zepdrix):

Oh... and the last term should be b^3 not d^3 I guess I spoke too soon :D lol

OpenStudy (love_333):

Oh crap, I messed up with my letters I was too focused on the numbers sorry

zepdrix (zepdrix):

XD

OpenStudy (love_333):

But it looks good as long as I correct those things?

zepdrix (zepdrix):

Yes.

OpenStudy (love_333):

Okay great! Thank you sooo much

zepdrix (zepdrix):

np c: